OpenStudy (anonymous):

The area of a regular hexagon is 38 cm^2. What is the area of a regular hexagon with sides eight times as long?

OpenStudy (anonymous):

area of a regular hexagon = \[\frac{ side^23\sqrt{3} }{ 2 }\] 38=side^2 3root/2 76=side^2 3root3 76=side^2root27 14.6262=side^2 side=3.82 so one side is 3.82 cm. now 3.82 times 8=? then just plug the known info back into the formula to solve. BTW my answer is rounded, so better if you do it yourself.

OpenStudy (accessdenied):

You might also see that because these figures have to be similar (they are both regular hexagons), there exists a ratio of their areas consistent with their ratio of side lengths 1:8. An area unit is just a length unit squared (A = k l^2), so the ratio between the areas is 1^2 : 8^2 or 1:64.

OpenStudy (anonymous):

Thanks you guys !

OpenStudy (accessdenied):

You're welcome. :)